15= 45/x <--------- Fraction.
Found 2 solutions by checkley77, rfer:
Answer by checkley77(12844)

15= 45/x
15x=45
x=45/15
x=3 ans.

Answer by rfer(16322)

15=45/x
multiply both sides by x
15x=45
divide both sides by 15
x=3